Minutes of the 10th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson

Minutes taker Mark Dennehy

Present Steven Watterson (Captain), Benjamin Jones (Treasurer), Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Conor McDermottroe (Public Relations Officer)

Absent(with apologies)

Absent Alan Smyth (Chief Range Officer), Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 26 September 2001 1800 – 2100

Distribution DURC Committee

Overview of topics

1 Correspondance

1 Correspondance 1 2 EARC 1 3 Freshers’ Week 2 4 Couch 2 5 Pony Club 2 6 Treasurers Report 2 7 New Range 3 8 Receipt books 3 9 NTSA Squad training 3 10 UCESSA 3 11 World Student Games 3 12 NTSA Squad 3 13 Beverage purchasing 4 14 Posters 4 15 DURC Squads 4 16 Freshers’ Week stand 4 17 Deirdre Ninaber 4 18 Irish Marksman 5

1 Correspondance

Conor received a letter from Pat Moray re the regulations for freshers’ week stands. It was read and discussed.

Mark had received notification of the UCESSA postal competitions.

2 EARC

2 EARC

Steven asked what the situation was regarding the EARC festival. Mark reported that the paperwork was uncertain, we had no ammunition and no confirmed logistics.

John reported talking to JK Walsh regarding ammunition. He (JK Walsh) will buy a bulk lot and we will purchase this lot from him in smaller (2-5000) lots. Delivery arrangements may be difficult but this was deemed a small price for such an arrangement.

Steven is to email Ray and Pete regarding travelling to EARC.

3 Freshers’ Week

John pointed out that we cannot accept last year’s ID badges as proof of studentship this year.

Steven suggested we wear our shooting jackets during freshers week as it would be more colourful. No slings however, due to the risk of injury in a crowd, and no glasses as they just look too wierd. Borrowing of jackets may occour, but borrowed jackets must not be fastened, they must be left open to prevent deforming. Mark also said that clean t-shirts must be worn underneath to limit soiling.

4 Couch

The broken couch is to be skipped before freshers week as it is unsafe and an eyesore. Ben’s GSU-donated chairs will suffice as replacements.

Steven moved the bin from beside the coffee machine to the range door.

5 Pony Club

Steven had spoken to Anne from the pony club regarding us coaching them. This prompted a long discussion on the relative merits of us investing in their coaching. Steven pointed out that they need different coaching techniques than we usually use. Anne wants a full day from us every three weeks on a weekend. This prompted complaints as weekend days are heavily subscribed allready. Ben pointed out that we see little or no short-term gain from this. Mark and Steven pointed out that it was good for the sport in the long term. It was voted on and agreed that we would coach them this year. A rota is to be drawn up and two people will make the trip every three weeks - thus noone should have to do this more than once every six to nine weeks. Conor and Peter Byrne can’t take part due to work committments.

DUEC were also mentioned. Mark wanted to know the status of our joint membership plan. Ben stated that we couldn’t afford to do this. Mark said that he thought the coaching was being done incorrectly, in that DUEC had one or two long sessions before their competition instead of lots of short ones. Steven wondered if we could shoot air rifle prone at 25yards for training. If so, we could have DUEC training in concert with normal prone training. John has a contact in DUEC - Laura O’Shea.

6 Treasurers Report

6 Treasurers Report

Ben reported sufficent funds for 3000 rounds of T22 and £120 to spend on freshers’ week’s reception. John pointed out that that was to be paid for from the monies recieved during Freshers’ week. Ben pointed out that the plan was to buy the supplies ahead of time from a cash’n’carry and thus upfront cash was required.

Mark asked if we had been lobbying for budget increases. Ben reported that we hadn’t really done so.

7 New Range

John pointed out that we had to start enquiring about the plans for the new range again from Terry McAuley. Ben reported new plans were up in DUCAC, but marked as not to scale. Tim Cooper told John that the Sports Complex committee’s minutes were now available on the web. Mark is to check on this.

8 Receipt books

All completed receipt books are to be returned to DUCAC.

9 NTSA Squad training

Mark reported the increase in charges for the national squad training weekends and the changes in procedures. Billing arrangements were discussed with Ben and it was agreed that the NTSA would notify us of the invitiation, we would reply with the number of people going, and they would send us a pro forma invoice. We would then arrive with a cheque.

10 UCESSA

Mark again noted the UCESSA postal entry forms. Entrys were solicited. For 25yd Prone, John, Mark, Conor and Peter Byrne will enter. For Air Rifle, Ben, Steven and Mark will enter. Other entries will be sought, the deadline being Oct.2.

Mark pointed out that due to the number of people who are disqualfied for not completing the course of fire each year, this year anyone disqualified for this reason will have to repay their entry fees to the club. This was agreed on by all.

11 World Student Games

Steven wondered what his status was in these games in regards to Nationality. Ben also wondered. Mark will try to find out.

12 NTSA Squad

12 NTSA Squad

Mark reported on the changes to the selection procedure for the national squads, and the formation of the developmental, junior, international and high performance squads. Entry levels were discussed, as were training arrangements. 2000 - At this

point John left

13 Beverage purchasing

Purchasing arrangements for supplies for the Freshers’ Week reception were discussed. Ben was asked for maximum amounts to spend and minimum amounts to obtain. £280/14 slabs of beer was agreed on. The question of quality versus quantity arose. It was decided that the bulk would be bought on a quantity basis, but if good quality beer was inexpensive enough, some would be purchased.

14 Posters

Conor showed his poster idea (posters printed on shot cards). These were agreed to be a good idea, which Conor attributed to Steven. It was decided to put them up in high-traffic areas in the Arts block and Engineering/Science areas.

15 DURC Squads

Mark asked for Steven’s thoughts on the DURC squads. Steven said that he thought the biggest problem was that we were not shooting regularly enough - three times a week being required to maintain a high standard, let alone progress. Two types of squad sessions were discussed - theory, which would be discussions of different aspects of shooting in a seminar format, and practical, which would entail shooting exercises and coaching. However, squad members would have to be training in between sessions by themselves as we cannot enforce this. Theory sessions might be a half-hour session every fortnight or so, with a topic discussed and then free-form discussion of anything people have problems with.

Entry to the squads should be determined by ladder scores. Squads should be of a managable size - no more than 10 per discipline.

16 Freshers’ Week stand

The stand must be set up and the rifles must be cleaned. People are too tired to do this at this point, so it was scheduled for Sunday afternoon as that was the next time most people could make it there. Steven pointed out that we are having too many committee meetings and not enough work sessions! Steven will procure duct tape and velcro for stand repairs.

Mark will start printing forms and leaflets and book the lecture hall.

17 Deirdre Ninaber

17 Deirdre Ninaber

Steven reported having spoken to Deirdre. She requires two weeks notice to get time from work, and has no computer access, which explains her absence through the summer.

18 Irish Marksman

Conor will write the DURC Report for the Irish Marksman as that was deemed to be the PRO’s job.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

18 Irish Marksman

Tasks agreed on at meeting